Giant Corporation uses standard costing. The following data are available for May: Actual costs of direct materials purchased $10,500 Actual price of direct materials $2.50 per kilogram Materials price variance $2,100 Favorable Materials spending variance $1,500 Favorable The quantity of materials purchased equals the quantity used in production. The standard quantity of material allowed for May production is
Answer #1

Material quantity variance = 1500-2100 = 600 U

Material price variance = (Standard price-actual price)actual quantity

2100 = (X*4200-10500)

-4200X = -12600

X(Standard price) = 3

Material quantity variance = (Standard qty-actual qty)Standard price

-600 = (X*3-4200*3)

-3X = -12000

X(Standard qty allowed) = 4000 kg